# On the volume of K-semistable Fano manifolds

### v1

@@ 2.3. Minimal rational curves
Since the celebrated work of Mori introducing the bend-and-break method for constructing rational curves on Fano manifolds ([Mor79]), the theory of rational curves has been developed extensively and has found numerous applications in algebraic geometry, especially in the classification of Fano manifolds with special properties. Here we recall some basic knowledge of rational curves and refer to [Kol96] for detailed expositions.
@@
$f^{*}TX={\mathcal{O}}_{{\mathbb{P}}^{1}}(2)\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)^{d-2}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}^{\oplus(n-d+1)}.$
Such a morphism $f$ must be an immersion and its image $f({\mathbb{P}}^{1})$ has at worst nodal singularities. We will also be interested in its normal bundle:
@@
Any Fano manifold always admits minimal rational curves, which can be obtained from the bend-and-break process starting with a free rational curve (see [Kol96, IV.Theorem 2.10]). Denote by $\mathrm{RatCurves}^{n}(X)$ the normalization of open subset of $\mathrm{Ch}(X)$ parametrizing integral rational curves. An irreducible component ${\mathcal{M}}$ of $\mathrm{RatCurves}^{n}(X)$ is referred to as a family of rational curves on $X$. The anticanonical degree $\deg({\mathcal{M}})$ of the family ${\mathcal{M}}$ is defined to be $-K_{X}\cdot C$ for any curve $C$ belonging to the family. This family ${\mathcal{M}}$ is equipped with a ${\mathbb{P}}^{1}$-bundle $p:{\mathcal{U}}\rightarrow{\mathcal{M}}$ and an evaluation morphism $q:{\mathcal{U}}\rightarrow X$. The family ${\mathcal{M}}$ is a dominating family if the evaluation morphism $q:{\mathcal{U}}\rightarrow X$ is dominant (i.e. has a dense image). This is equivalent to the condition that a general rational curve in ${\mathcal{M}}$ is free. A dominating family ${\mathcal{M}}$ is locally unsplit if, for a general point $x\in X$, the subfamily ${\mathcal{M}}_{x}=p(q^{-1}(x))$ parametrizing curves through $x$ is proper (i.e. compact). Note that ${\mathcal{M}}$ is a dominating family of rational curves on $X$ that has a minimal anticanonical degree, then ${\mathcal{M}}$ is a dominating locally unsplit family. But not all dominating locally unsplit family has a minimal anticanonical degree.
Let ${\mathcal{M}}$ be a dominating locally unsplit family of rational curves. By Mori’s bend-and-break argument, we know that for a general point $x\in X$, a general curve in ${\mathcal{M}}_{x}$ is minimal. In particular, $\deg({\mathcal{M}})\in\{2,3,\cdots,n+1\}$. Indeed, if a general free rational curve $C=[f]$ in ${\mathcal{M}}_{x}$ is not minimal, then there are at least two ${\mathcal{O}}(2)$ summands in the splitting of $f^{*}TX$. We can then fix two points on the curve $C$ and bend-and-break rational curve into a non-integral (reducible) curve which would contradict the locally unsplit property. Following [Miy04, CD15], we set
(6) $l_{X}:={\rm min}\{\deg{\mathcal{M}};{\mathcal{M}}\text{ is a dominating locally unsplit family of rational curves on }X\}.$
By the above discussion, it is easy to see that (see [CD15, Remark 4.2]):
@@
Example 2.9.
To prove our main results, we will need the following important results on classification of Fano manifolds with dominating locally unsplit families of large anticanonical degrees.
@@
    ([Miy04, CD15, DH17]) $l_{X}=n$ if and only if $X\cong Q^{n}$ or $X$ is the blowup of ${\mathbb{P}}^{n}$ along subvariety $Y$ of degree $d_{Y}\in\{1,\dots,n\}$ that is contained in a hyperplane.
@@
    for a general point $x\in X$, there exists a rational curve which has degree 2 with respect to $K_{X}^{-1}$.
@@
3. Warm-up: case of trivial normal bundles
In this section, we prove the following theorem, which is a special case of Theorem 1.1.
Theorem 3.1.
Let $X$ be a K-semistable Fano manifold that contains a rational curve with a trivial normal bundle. Then $(-K_{X})^{n}\leqslant 2n^{n}$ and the equality holds if and only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
The rest of this section is devoted to the proof of Theorem 3.1. We start with a proposition that proves a more general version of estimate (1) by incorporating the $\delta$-invariant.
@@ 3. Warm-up: case of trivial normal bundles
Assume $X$ is a $n$-dimensional Fano manifold, and $Z\subset X$ is a codimension $r$ non-singular subvariety with trivial normal bundle $N_{Z/X}={\mathcal{O}}_{Z}^{\oplus r}$, and we set $d=(-K_{X})^{n-r}\cdot Z$, then
@@
$0\rightarrow{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j}\rightarrow 0$
@@
The argument is similar with [Fuj18, Theorem 2.3(2)]. We denote
@@
$0\rightarrow H^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})\rightarrow H^{0}(X,kL)\rightarrow H^{0}(X,kL\otimes{\mathcal{O}}_{kx_{t}Z})\rightarrow H^{1}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})\rightarrow 0$
@@
$0\rightarrow{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j}\rightarrow 0$
@@
When $X$ is a smooth toric Fano manifold, by the work of Chen-Fu-Hwang [CFH14, Corollary 2.5] which is partly based on the work of Araujo [Ara06], we know that there exists submanifold $Z\cong{\mathbb{P}}^{n-r}$ that is contained in $X$ with trivial normal bundle. By Proposition 3.2 and $\delta(X)\geqslant 1$, we have $(-K_{X})^{n}\leqslant(-K_{{\mathbb{P}}^{r}\times Z})^{n}=(-K_{{\mathbb{P}}^{r}\times{\mathbb{P}}^{n-r}})^{n}=\binom{n}{r}\cdot(r+1)^{r}\cdot(n-r+1)^{n-r}=:c_{r}$. By the following lemma 3.5, for $12\leq r\leq n-$ 1, $c_{r}\leq 2n^{n}$ with equality holds if and only if $r=2$ or $r=n-1$ in which case $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$. ∎
@@
[…]
@@
Let $X$ be a K-semistable Fano manifold. Assume that $f:{\mathbb{P}}^{1}\rightarrow X$ is an immersed rational curve with a non-empty subset of nodal points such that $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}^{n-1}$. Then we have an estimate of the volume $(-K_{X})^{n}<2n^{n}$. As a consequence, if there exists a minimal rational curve with a trivial normal bundle on $X$ and $(-K_{X})^{n}=2n^{n}$, then the minimal rational curve has no nodal points and $X$ must be isomorphic to ${\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
Proof.
The last statement follows from the strict volume inequality and Theorem 1.5. So we just need to prove the inequality $(-K_{X})^{n}<2n^{n}$. Denote the ideal sheaf of the nodal curve $C=f({\mathbb{P}}^{1})$ in $X$ by ${\mathcal{I}}={\mathcal{I}}_{C}$. Consider the blow up $\mu:\tilde{X}\rightarrow X$ along $C$. By using the same argument as in the proof of Proposition 3.2, we get the estimate:
${\rm vol}({\mu}^{*}(-K_{X})-xE)$$\geqslant$ $(-K_{X})^{n}-\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{h^{0}(X,L^{k}\otimes{\mathcal{O}}_{kxC})}{k^{n}/n!}$
$\geqslant$ $(-K_{X})^{n}-\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{\mathop{\sum}_{y=0}^{xk-1}h^{0}(C,kL\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})}{k^{n}/n!}.$
Since $C$ is a local complete intersection, its conormal sheaf $N^{\vee}_{C/X}={\mathcal{I}}/{\mathcal{I}}^{2}$ is locally free and ${\mathcal{I}}^{y}/{\mathcal{I}}^{y+1}={\rm Sym}^{y}(N_{C/X}^{\vee})$. So we get:
$h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})=h^{0}(C,L^{k}\otimes{\rm Sym}^{y}({\mathcal{I}}/{\mathcal{I}}^{2}))\leqslant h^{0}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}(dk)\otimes{\rm Sym}^{y}(f^{*}({\mathcal{I}}/{\mathcal{I}}^{2}))).$
There is an inclusion ${\mathcal{F}}:=f^{*}({\mathcal{I}}/{\mathcal{I}}^{2})\hookrightarrow N_{f/X}^{\vee}=(f^{*}TX/T{\mathbb{P}}^{1})^{\vee}$. We know that $\deg(N^{\vee}_{f/X})=2-d$ and $\deg({\mathcal{F}})=\deg(K_{X})-\deg\omega_{C}=-d-2m+2$ where $d=-K_{X}\cdot C$. Since $N^{\vee}_{f/X}$ is trivial, $d=2$ and ${\mathcal{F}}={\mathcal{O}}(-a_{1})\oplus{\mathcal{O}}(-a_{2})\oplus\cdots\oplus{\mathcal{O}}(-a_{n-1})$ with $a_{1}\geq a_{2}\geq\cdots\geq a_{n-1}\geq 0$ and $\sum_{i=1}^{n-1}a_{i}=2m$. It is then easy to get the estimate:
(12) $h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leqslant h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(2k)\otimes{\rm Sym}^{y}({\mathcal{O}}(-2)\oplus{\mathcal{O}}^{\oplus(n-2)}))$
or
(13) $h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leqslant h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(2k)\otimes{\rm Sym}^{y}({\mathcal{O}}(-1)^{\oplus 2}\oplus{\mathcal{O}}^{\oplus(n-3)})).$
For the first case (12), we get:
(14) $\sum_{y=0}^{xk-1}h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leq\sum_{y=0}^{xk-1}\sum_{i=0}^{y}(2k-2i+1)a_{y,i}$
where $a_{y,i}=\binom{n-3+y-i}{n-3}$. If $x\leqslant 1$, then by using the explicit expression of binomial coefficient, the right-hand-side of (14) is equal to $b_{1}k^{n}/n!+O(k^{n-1})$ where $b_{1}$ is given by the following integral (see (19) for a derivation of a more complicated formula):
$b_{1}$ $=$ $n!\int_{0}^{x}{\rm d}t\int_{0}^{t}(2-2z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z=\frac{2n!}{(n-2)!}\int_{0}^{x}(1-z)(x-z)^{n-2}{\rm d}z$
$=$ $2nx^{n-1}-2x^{n}<2nx^{n-1}.$
When $x>1$, the right-hand-side of (14) splits into the sum: $\sum_{y=0}^{k}\sum_{i=0}^{y}+\sum_{y=k+1}^{xk-1}\sum_{i=0}^{k}$ and the right-hand-side of (14) is equal to $b_{2}k^{n}/n!+O(k^{n-1})$ with
$b_{2}$ $=$ $(2n-2)+n!\int_{1}^{x}{\rm d}t\int_{0}^{1}(2-2z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z$
$=$ $(2n-2)+\frac{2n!}{(n-2)!}\int_{0}^{1}(1-z)((x-z)^{n-2}-(1-z)^{n-2}){\rm d}z$
$=$ $\frac{2n!}{(n-2)!}\int_{0}^{1}u(x-1+u)^{n-2}{\rm d}u+(2n-2)-(2n-2)$
$=$ $2\displaystyle 2(nx^{n-1}-x^{n}+(x-1)^{n})<2(n-1)x^{n-1}+$ 2 .
We set $\psi=2nx^{n-1}$ and
(15) ${\phi}(x)=\left\{\begin{array}[]{ll}2nx^{n-1}&\text{if }\;0\leq x\leq 1\\ 2(n-1)x^{n-1}+2&\text{if }\;x>1,\end{array}\right.$
so that $\phi(x)<\psi(x)$. It is easy to check that $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ and the strict inequality holds when $V>2n$. So we can apply Lemma 2.4 to conclude that $F^{\phi}(V)\geq F^{\psi}(V)$ and the strict inequality holds for $V>2n$. From Lemma 2.4 and Corollary 2.5, we get $(-K_{X})^{n}\leq(F^{\phi})^{-1}(n-1)<(F^{\psi})^{-1}(n-1)=2n^{n}$.
In the second case (13), we get:
$\sum_{y=0}^{xk-1}h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leq\sum_{y=0}^{xk-1}\sum_{i=0}^{y}(2k-i+1)(i+1)\binom{n-4+y-i}{n-4}.$
Estimating as before, when $x\leqslant 2$, the right-hand-side of (14) is equal to $c_{1}k^{n}/n!+O(k^{n-1})$. The coefficient $c_{1}$ is equal to:
$c_{1}$ $=$ $n!\int_{0}^{x}{\rm d}t\int_{0}^{t}(2-z)z\frac{(t-z)^{n-4}}{(n-4)!}{\rm d}z$
$=$ $\frac{n!}{(n-3)!}\int_{0}^{x}(2-z)z(x-z)^{n-3}{\rm d}z=2x^{n-1}(n-x)<2nx^{n-1}.$
When $x\geqslant 2$, the right-hand-side of (14) splits into the sum: $\sum_{y=0}^{2k}\sum_{i=0}^{y}+\sum_{y=2k+1}^{xk-1}\sum_{i=0}^{2k}$ and the right-hand-side of (14) is equal to $c_{2}k^{n}/n!+O(k^{n-1})$ with the coefficient $c_{2}$ is equal to:
$c_{2}$ $=$ $2^{n}(n-2)+n!\int_{2}^{x}{\rm d}t\int_{0}^{2}(2-z)z\frac{(t-z)^{n-4}}{(n-4)!}{\rm d}z$
$=$ $2^{n}(n-2)+\frac{n!}{(n-3)!}\int_{0}^{2}(2-z)z((x-z)^{n-3}-(2-z)^{n-3}){\rm d}z$
$=$ $\frac{n!}{(n-3)!}\int_{0}^{2}(2-z)z(x-z)^{n-3}{\rm d}z$
$=$ $2nx^{n-1}-2x^{n}+(x-2)^{n}+n(x-2)^{n-1}<(2n-1)x^{n-1}$
$<$ $(2n-1)x^{n-1}+2^{n-1}.$
We set $\psi(x)=2nx^{n-1}$ and
(16) ${\phi}(x)=\left\{\begin{array}[]{ll}2nx^{n-1}&\text{if }\;0\leq x\leq 2\\ (2n-1)x^{n-1}+2^{n-1}&\text{if }\;x\geq 2.\end{array}\right.$
It is easy to check that $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ and strict inequality holds for $V>2^{n}\cdot n$. Then by Lemma 2.4 we get $F^{\phi}(V)\geq F^{\psi}(V)$ with strict inequality when $V>2^{n}\cdot n$. So we get $(-K_{X})^{n}\leq(F^{\phi})^{-1}(n-1)<(F^{\psi})^{-1}(n-1)=2n^{n}$. ∎
Proof of Theorem 3.1.
If $X$ contains a rational curve with non-empty nodal points, then by Proposition 3.6, we have $(-K_{X})^{n}<2n^{n}$. So it suffices to consider the case when $C$ is an embedded rational curve, in which case the result follows immediately from Theorem 1.5. ∎
Proposition 3.7.
@@
Suppose $X$ is a singular $n$-dimensional K-semistable toric ${\mathbb{Q}}$-Fano variety. When $n=2$, then $(-K_{X})^{2}\leqslant\frac{9}{2}$. When $n\geqslant 3$, then $(-K_{X})^{n}\leqslant\frac{16}{27}(n+1)^{n}$. In particular, we have $(-K_{X})^{n}<(-K_{{\mathbb{P}}^{1}\times{\mathbb{P}}^{n}})^{n}=2n^{n}$ holds for all positive integer $n\geqslant 2$.
@@
The second largest volume of $n$-dimensional K-semistable toric ${\mathbb{Q}}$-Fano varieties is $2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
@@
We conclude this section by a convex-geometric application.
@@
Suppose $P\subseteq{\mathbb{R}}^{n}$ is a $n$-dimensional reflexive lattice polytope with barycenter at $0\in{\mathbb{R}}^{n}$. Assume $P$ is not unimodularly equivalent to $(n+1)$ times a standard simplex $(n+1)\Delta_{n}$, then the volume of $P$ with respect to the Lebesgue measure in ${\mathbb{R}}^{n}$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and the equality holds if and only if $P$ is unimodularly equivalent to $[0,2]\times(n\Delta_{n-1})$.

### v2

@@ 2.3. Minimal rational curves
Since the celebrated work of Mori that introduces the bend-and-break method for constructing rational curves on Fano manifolds ([Mor79]), the theory of rational curves has been developed extensively and has found numerous applications in algebraic geometry, especially in the classification of Fano manifolds with special properties. Here we recall some basic knowledge of rational curves and refer to [Kol96] for detailed expositions.
@@
$f^{*}TX={\mathcal{O}}_{{\mathbb{P}}^{1}}(2)\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}^{\oplus(n-d+1)}.$
Such a morphism $f$ must be an immersion. We will also be interested in its normal bundle:
@@
Any Fano manifold always admits minimal rational curves, which can be obtained from the bend-and-break process starting with a free rational curve (see [Kol96, IV.Theorem 2.10]). Denote by $\mathrm{RatCurves}^{n}(X)$ the normalization of open subset of $\mathrm{Ch}(X)$ parametrizing integral rational curves. An irreducible component $\mathcal{M}$} of $\mathrm{RatCurves}^{n}(X)$ is referred to as a family of rational curves on $X$. The anticanonical degree $\deg({\mathcal{M}})$ of the family $\mathcal{M}$} is defined to be $K_{X}\cdot C$ for any curve $C$ belonging to the family. This family $\mathcal{M}$} is equipped with a ${\mathbb{P}}^{1}$-bundle $p:{\mathcal{U}}\rightarrow{\mathcal{M}}$ and an evaluation morphism $q:{\mathcal{U}}\rightarrow X$. The family $\mathcal{M}$} is a dominating family if the evaluation morphism $q:{\mathcal{U}}\rightarrow X$ is dominant (i.e. has a dense image). This is equivalent to the condition that a general rational curve in $\mathcal{M}$} is free. A dominating family $\mathcal{M}$} is locally unsplit if, for a general point $x\in X$, the subfamily ${\mathcal{M}}_{x}=p(q^{-1}(x))$ parametrizing curves through $x$ is proper (i.e. compact). Note that if $\mathcal{M}$} is a dominating family of rational curves on $X$ that has a minimal anticanonical degree, then $\mathcal{M}$} is a dominating locally unsplit family. But not all dominating locally unsplit family has a minimal anticanonical degree.
Let $\mathcal{M}$} be a dominating locally unsplit family of rational curves. By Mori’s bend-and-break argument, we know that for a general point $x\in X$, a general curve in ${\mathcal{M}}_{x}$ is minimal. In particular, $\deg({\mathcal{M}})\in\{2,3,\cdots,n+1\}$. Indeed, if a general free rational curve $C=[f]$ in ${\mathcal{M}}_{x}$ is not minimal, then there are at least two ${\mathcal{O}}(2)$ summands in the splitting of $f^{*}TX$. We can then fix two points on the curve $C$ and bend-and-break the rational curve into a non-integral (reducible) curve which would contradict the locally unsplit property. Following [Miy04, CD15], we set
[…]
By the above discussion, we see that (see [CD15, Remark 4.2]):
@@
Theorem 2.9.
Using the above notation, we have the following important classificaiton results:
@@
    ([Miy04, CD15, DH17]) $l_{X}=n$ if and only if $X\cong Q^{n}$ or $X$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y$ of degree $d_{Y}\in\{1,\dots,n\}$ that is contained in a hyperplane.
@@
    for a general point $x\in X$, there exists a rational curve that has degree 2 with respect to $K_{X}^{-1}$.
@@
3. Submanifolds with normal bundles
In this section, we will prove Theorem 1.4 by generalizing the methods from [Fuj18, LZ18], and apply it in the toric case to prove Theorem 1.2. We start with a proposition that proves a more general version of the estimate (1) by incorporating the $\delta$-invariant.
@@ 3. Submanifolds with normal bundles
Let $X$ be a $n$-dimensional Fano manifold, and $Z\subset X$ be a codimension $r$ non-singular subvariety with trivial normal bundle $N_{Z/X}={\mathcal{O}}_{Z}^{\oplus r}$. Set $d=(-K_{X})^{n-r}\cdot Z$. Then we have the estimate:
@@
$0\rightarrow{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j}\rightarrow 0.$
@@
The argument is similar to the proof of [Fuj18, Theorem 2.3(2)]. We denote
@@
$0\rightarrow H^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})\rightarrow H^{0}(X,kL)\rightarrow H^{0}(X,kL\otimes{\mathcal{O}}_{kx_{t}Z})\rightarrow H^{1}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})\rightarrow 0.$
@@
$0\rightarrow{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j}\rightarrow 0.$
@@
When $X$ is a smooth toric Fano manifold, by the work of Chen-Fu-Hwang [CFH14, Corollary 2.5] which is partly based on the work of Araujo [Ara06], we know that there exists a submanifold $Z\cong{\mathbb{P}}^{n-r}$ for $0\leq r\leq n-1$ that is contained in $X$ with a trivial normal bundle. By Proposition 3.1 and $\delta(X)\geqslant 1$, we have $(-K_{X})^{n}\leqslant(-K_{{\mathbb{P}}^{r}\times Z})^{n}=(-K_{{\mathbb{P}}^{r}\times{\mathbb{P}}^{n-r}})^{n}=\binom{n}{r}\cdot(r+1)^{r}\cdot(n-r+1)^{n-r}=:c_{r}$. By the following lemma 3.4, for $1\leq r\leq n-1$, $c_{r}\leq 2n^{n}$ with equality holds if and only if $r=1$ or r=n1r=n-1 and in both cases $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$. ∎
@@
where $a_{k}=(\frac{k+1}{k})^{k}=(1+k^{-1})^{k}$ is an strict increasing sequence for $k\geq 1$ (that converges to $e$ as $k\rightarrow+\infty$). So we get $c_{r+1}/{c_{r}}<1$ when r+1<nrr+1<n-r (or equivalently $2r+1<n)$. The statement then follows easily. ∎
@@
Let $X$ be a singular $n$-dimensional K-semistable toric $\mathbb{Q}$}-Fano variety. When $n=2$, then $(-K_{X})^{2}\leqslant\frac{9}{2}$. When $n\geqslant 3$, then $(-K_{X})^{n}\leqslant\frac{16}{27}(n+1)^{n}$. In particular, we have $(-K_{X})^{n}<(-K_{{\mathbb{P}}^{1}\times{\mathbb{P}}^{n}})^{n}=2n^{n}$ holds for all positive integer $n\geqslant 2$.
@@
The second largest volume of $n$-dimensional K-semistable toric $\mathbb{Q}$}-Fano varieties is $2n^{n}$ and it is obtained only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
@@
As mentioned in the Introduction, Theorem 3.6 finds an application in Arakelov geometry on the sharp bound of height of arithmetic Fano varieties whose complexification is K-semistable. We refer our readers to [AB24, Ber25] for more detailed explanation and possible arithmetic applications.
Theorem 3.8.
Andreasson-Berman’s conjecture holds for the toric canonical model of a toric Fano varieties. Namely, for a $n$-dimensional K-semistable toric Fano variety and let $\mathcal{X}$} be the canonical model of $X$ over $\mathbb{Z}$}, then for any volume-normalized continous metric $\left\|\cdot\right\|$ on $K_{X}$ with positive curvature current, we have $h_{\left\|\cdot\right\|}({\mathcal{X}},-K_{{\mathcal{X}}})\leq h_{\left\|\cdot\right\|_{\text{FS}}}({\mathbb{P}}^{n}_{{\mathbb{Z}}},-K_{{\mathbb{P}}^{n}_{{\mathbb{Z}}}})$ where $K_{{\mathbb{P}}^{n}_{{\mathbb{C}}}}$ is endowed with the volume normalized Fubini-Study metric.
Proof.
It follows directly from Theorem 3.6 and [AB24, Lemma 3.8]. ∎
We conclude this section with another application in convex geometry.
@@
Suppose $P\subseteq{\mathbb{R}}^{n}$ is a $n$-dimensional reflexive lattice polytope with its barycenter at $0\in{\mathbb{R}}^{n}$. Assume that $P$ is not unimodularly equivalent to $(n+1)$ times a standard simplex $(n+1)\Delta_{n}$, then the volume of $P$ with respect to the Lebesgue measure in ${\mathbb{R}}^{n}$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and the equality holds if and only if $P$ is unimodularly equivalent to $0,2]\times(n\Delta_{n-1})$.